OpenStudy (anonymous):

A dice is rolled 6 times. Let S be the sum, over the 6 rolls, of the number on the face that turned up. The probability that S ≥ 35 is: . (Write your answer in the form 0.xxxxxx, rounded up to the 6th digit.)

OpenStudy (anonymous):

this can be done like find the coeffecient of x^35 in (x+x^2+x^3+x^4+x^5+x^6)^6 ... then divide this by 6^6 you would get the probability ... you have to apply binomial theorem

OpenStudy (anonymous):

=6/6^6 = 1/6^5 .. check the calculation .. I did it in excel.. but do check once

OpenStudy (anonymous):

is it 0.000150?